Depreciation Expense Using the Double-Declining Balance Method
The Pack Company purchased an office building for $4,500,000. The building had an estimated useful life of 40 years and an expected salvage value of $500,000.

Calculate the depreciation expense for the second year using the double-declining balance method. $Answer is 213750

can you plz tell me how can I get this number ?
Answer #1

Double declining rate = 200/useful life = 200/40 = 5% per year

Depreciation for first year

= 4,500,000 * 5% = 225,000

Depreciation for second year

= (4,500,000-225,000) * 5% = 213,750